# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v1

@@ III-A Attacks
where $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$. Given its dimensionality, it can be intractable to learn such a policy that modifies the input/problem space directly [39]; CIFAR-10, for instance, has more than 3K features to perturb. In AI enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can assume that $F_{0}=F_{1}=...=F_{t},\forall t$. While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand. The existence of adversaries which follow a candidate generation policy introduces however a behavior which can be observed and utilized by a defensive methodology. Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending against decision-based attacks.
@@
Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$. Then given an HSJA attack, in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient. If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to decision-based attacks which navigate the boundary. Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
@@ III-B Defenses
Proposition III.1 shows that alternative classification policies are useful in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection. In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [7, 18]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [24, 44]. While adversarially training the discriminant function $F$ empirically shows the capacity to resist decision-based attacks, the manner in which the model responds has a complementary potential. This gap between the empirical and theoretically possible robustness to decision-based attacks is the locus where a distinct from model hardening, active and adaptive defense can emerge. Active defenses have direct implications on attacks themselves however. Let us assume an agent carrying out an active defense policy:
$\pi_{\phi}^{\mathcal{D}}=P(\alpha|x_{t},s^{\mathcal{D}}),\>a\in\{0,1\}$ (7)
with $x_{t}$ the query, $s^{\mathcal{D}}$ the state for the defense as created by past queries, and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$. If this policy is stationary the environment dynamics become in turn stationary and so, next to the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved. In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [49]. Active defenses, as a consequence of decision-based attacks, entail adaptive adversaries.
@@
A proof for BAGS and HSJA is included in Appendix A. An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too. However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often not optimal out-of-the-box. Adapting attack policies can be seen as performing hyperparameter optimization and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [57, 38]. The empirical results in Section V further indicate this correspondence between adaptive and self-optimizing, where the adaptive versions of attacks outperform the non-adaptive, more so against active defenses.
@@
Consider now an active defense that is based on a similarity or a conformal metric. In the twofold meaning of adaptive we introduced in Section II, inventing control implies the capability to bypass a similarity based defense; adaptive control implies strategy instead, the active control of the available tools to evade the defense [1]. Then the adversary’s task is to find the optimal adversarial policy that also evades rejection. The straightforward way to achieve this is by adapting policy (2) itself. Notably, and despite the discrete and black-box context, this optimization can be fully gradient-based [48]. We now demonstrate how adaptively controlling decision-based attacks recovers the gradient-based solvability of the black-box optimization task despite neither the active defense nor the model itself being accessible in closed-form.
Theorem III.1 (Adversarial Policy Gradient).
@@
Given model $\mathcal{M}$ with an active and stationary defense $\pi_{\phi}^{\mathcal{D}}$ (7), adaptive candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ (2) that generates episodes $\tau$ of queries $x_{t}$, and a reward $r(\tau)=\sum_{x_{t}\in\tau}\neg\alpha$, the optimal evasive policy $\mathcal{E}$ is obtained via the gradient of the policy’s expected reward $\mathbb{E}_{\pi^{\mathcal{A}}_{\theta}}[r(\tau)]$.
@@
The proof is included in Appendix A. So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) by mere observation of discrete model decisions, adaptive attacks can become optimal in terms of both evasion and efficiency. The last piece of the puzzle is turning active defenses also adaptive.
Proposition III.3.
@@
An active defense $\pi_{\phi}^{\mathcal{D}}$ (7) achieves its optimal, i.e. maximizes the expectation $\mathbb{E}[\sum_{x_{t}\in\tau}P(\alpha|x_{t},s)]$, by adapting its policy against a stationary evasive policy $\mathcal{E}$.
@@
As offensive and defensive policies are strictly competitive, we can define the reward $\rho$ of the defensive policy as $\rho(\tau)=\sum_{x_{t}\in\tau}\alpha$, then by making $\pi_{\theta}^{\mathcal{A}}$ stationary and $\pi_{\phi}^{\mathcal{D}}$ adaptive in Theorem III.1, we can reason that the optimal defensive policy is determined via the gradient of the expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau)]$. ∎
Figure 2: Schematic model of an AMG environment. Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations.
@@ III-C Adversarial Markov Games
By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently from each other. As adaptive decision-based attacks and defenses are logical consequences of each other, by composing them we can form a turn-taking competitive game. A precise game-theoretic formulation requires the exact analytical description of the whole environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this context as well as most cybersecurity environments. Model-free methods however can learn optimal offensive and defensive responses directly through interaction with the environment [44, 43], obviating the need to learn a model of it or to find exact solutions the to bi-level optimization task like Eq. (1) that is inherently NP-hard to solve [11].
To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al. [25] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information; an apt formalism for decision-based attacks and defenses. Another nice property of TT-POMGs is that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
The competition underlying adversarial example generation has been explored in no-box and white-box settings [8, 22]. We instead focus on decision-based, interactive environments which are assumed to have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior. By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG. The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
This environment that encompasses adversarial attacks, adversarial defenses, and benign queries, can be construed as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2. Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and as a consequence the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be adaptive in turn. As without implausible assumptions one cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game. When human agents compete by holding beliefs about each other, they engage in recursive reasoning that in theory of mind is encountered as [I believe that [my opponent believes [that I believe…]]]. In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [2, 54]. However, as AMGs can be solved by single-agent RL algorithms, we consider more involved recursive reasoning out of scope and perform the empirical evaluation without building explicit models of opponent behavior.
@@ IV Environment Specification
The empirical study we conduct in Section V constitutes diverse instantiations of the general theoretical framework introduced in Section III. Naturally, when working forward from the general to the particular concrete design choices have to be made when specifying the latter, and these choices can have considerable influence on the results. In this section we thus outline the components of the environment: adversaries, defenders, possible interactions and various mechanisms that underscore them.
@@
While extensible to more players, our AMG framework describes a two-player competitive game, so we assume that at a given moment only one attack takes place. From the defensive perspective, incoming queries can be either benign (clean) or part of an attack. A frequently made assumption that influences the effectiveness of stateful detection approaches, however, is the ability to attribute queries to UIDs, e.g., an IP address or a user account. Adversaries can create multiple accounts, use VPNs, collude between them, or in fact accounts or IP addresses might not even be necessary to query the model. To address that, we treat queries irrespective to their source, a strictly more challenging setting for stateful defenses. The environments we build operate solely on the content of queries and not on any other metadata, similar to [32]. Like this work, we reframe the task of detecting and banning UIDs of adversaries with two notable differences. Instead of rejecting queries, a decision that in itself can assist the adversary in optimizing, we misdirect by returning another decision, e.g. the second highest class. Secondly, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution so that is not actually trivial to tell them apart from adversarial noise.
@@
Similarity. As decision-based attacks are governed by a generation policy, successive queries exhibit patterns and degrees of similarity that can be quantified by an appropriate $l_{p}$ norm. However such a norm should not be computed on the original inputs as an adapted adversary can ”invent control” (see Fig. 1) by employing evasive transformations the model is invariant to and bypass the similarity detection. To account for such adversarial strategies, we train a Siamese network with contrastive loss in order to learn an latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs. To statefully characterize queries we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign/undetected queries.
@@
Figure 3: Intuition of misdirection in a two-dimensional decision boundary. The adaptive defense controls a single parameter: the radius around $c_{0}$ (the last known adversarial) where queries $x_{t}$ side are intercepted and the model responds with a non-adversarial decision. $x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
@@
Adaptivity. No evaluation in AML is complete without considering adaptive adversaries: that is with the instruments to bypass the defense, but not necessarily with their optimal configuration. As all stateful defenses so far are similarity based, to bypass them our intuition points towards input transformations the model is invariant to. Formally, for a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))=F(x_{t})$. Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold. As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
@@
So what is the correct composition of transformations $T$ to apply? When shall $T$ be applied, and how does it affect the attack fundamentals? The possible transformations can be considered as set of additional controls, and like attack parameters they themselves can be underperforming out-of-the-box [19]. Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense: their empirical configuration is often suboptimal against a particular defense and a fine balance is often required between evasion and preserving attack performance. This trade-off illustrates why the twofold definition of adaptive we trace in this work is necessary in AML evaluations: first to impart controllability to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
@@
Agents & Environments. Unlike the vast majority of competitive games, in AMGs the two players have different action and state sets. AMGs are also asymmetric in the playing cadence: while the defender plays every round, it takes one to several rounds for the adversary to play; HSJA for example is controlled on the iteration rather on the query level. The learning task is complicated further given that the experience upon which each agent learns arrives only after the opponent moves. We address these complications by developing custom learning environments where agents can handle asymmetries and the delayed experience collection by training asynchronously. For the environments and agents respectively, we build on the OpenAI Gym and Stable Baselines 3 libraries  [10, 41]. The exact definitions of the RL states and rewards we experimented with are included in Appendix B.
@@
States & Actions. For the definition of the states we used and the rationale behind them, we point the reader to Appendix B. For actions, we control BAGS through 4 parameters: orthogonal step size, source step size, mask bias, and Perlin bias. HSJA is controlled by 3: the gradient estimation radius, the number of estimation queries, and the jump step size. All evaluations start from controlling these fundamental attack parameters only. If the active defense proves impossible to defeat, we introduce additional knobs corresponding to the evasive actions that control the magnitude and probability of transformations on the input, with the goal to evade detection while preserving semantic content and hence the correct classification. The range of transformations we experimented with as well as their magnitude and probability are listed in Table V. Finally, in both BAGS and HSJA the active defense corresponds to an 1-dimensional continuous action that controls the radius of confinement, as depicted in Fig 3.
@@
Rewards. The success of any RL task relies heavily on how it is rewarded. Engineering an effective reward function is often non-trivial and can contain many intricacies, as reward hacking and specification gaming are common phenomena and the learned behavior can vary significantly [3]. For adversaries, the rewards we experimented with are variations on minimizing the distance to the original example, while defenders are rewarded and penalized for intercepting adversarial and benign queries respectively. More specifically, BAGS terms include tn the distance gains towards the original example (also relative to the initial gap) and on their frequency. For HSJA, additional terms include penalties for the amount of gradient estimation queries, and rewards on the precision of gradient estimation and jump steps. Defensive rewards for BAGS include keeping the best adversarial or the average query close to the starting point, minimizing the size of the step, and correctly intercepting queries. For HSJA, additional terms include penalties for not intercepting queries at the binary search or jump steps. The reward functions in closed-form are shown in Appendix B.

### v2

@@ III-A Attacks
where $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$. Given the dimensionality, it can be intractable to learn a policy that modifies the inputs at the feature space directly [36]; CIFAR-10, for instance, has more than 3K features to perturb.
In AI-enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$. While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand. The fact however that adversaries follow a candidate generation policy, introduces a querying behavior which can be observed and exploited by a defensive methodology. Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending from decision-based attacks.
@@
Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$. Then in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$ in HSJA, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient. If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to other decision-based attacks which navigate the boundary. Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
@@ III-B Defenses
Proposition III.1 indicates that alternative classification policies are necessary in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection. In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [11, 37]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12]. While adversarially training the discriminant function $F$ empirically shows the capacity to resist decision-based attacks, the manner in which the model responds has a complementary potential. This gap between the empirical and theoretically possible robustness to decision-based attacks is the locus where a distinct from model hardening, active and adaptive defense can emerge. Active defenses have direct implications on attacks themselves however. Let us now assume an agent carrying out an active defense policy:
$\pi_{\phi}^{\mathcal{D}}=P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t}),\>\alpha\in\{0,1\}$ (7)
with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$. If this policy is stationary, the environment dynamics in turn become stationary too, and thus together with the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved. In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [38]. Active defenses, a consequence of decision-based attacks, entail therefore adaptive adversaries.
@@
A proof for BAGS and HSJA is included in Appendix A. An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too. However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often suboptimal with their default, empirically defined parameters. Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [39, 40, 41]. Our results in Section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive attacks consistently outperform non-adaptive, particularly against active defenses.
@@
Consider now an active defense that is based on a similarity or a conformal metric. In the twofold meaning we introduced in Section II, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf. Fig. 1). The revised adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself. Notably, and despite the black-box and discontinuous nature of the task, this optimization can be fully gradient-based. Decision-based attacks can recover gradient-based solutions to their objective, despite neither the active defense nor the model itself being accessible in closed-form. For model $\mathcal{M}$, adversarial queries $x_{t}$, and active defense $\pi_{\phi}^{\mathcal{D}}$ making decisions $\alpha_{t}$, we can thus formulate the following:
Proposition III.3 (Adversarial Policy Gradient).
@@
Given adversarial policy $\pi_{\theta}^{\mathcal{A}}$ (2) that generates episodes $\tau_{i}$ of queries $x_{t}$, and reward function $r(\tau_{i})=\sum_{x_{t}\in\tau_{i}}(1-\alpha_{t})$, the optimal evasive policy ${\pi^{\mathcal{A}}_{\theta^{*}}}$ is obtained by gradient ascent on the policy’s expected reward, $\nabla_{\theta}\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]$.
@@
The proof is included in Appendix A. So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions. To complete the puzzle, the last piece is turning active defenses also adaptive.
Corollary III.4.
@@
The active defense achieves its optimal $\pi_{\phi^{*}}^{\mathcal{D}}$ (7), i.e. maximizing expectation $\mathbb{E}[\sum_{x_{t}\in\tau_{i}}P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t})]$, by adapting its policy against the optimal evasive policy ${\pi^{\mathcal{A}}_{\theta^{*}}}$.
Figure 2: Schematic model of an AMG environment. Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations. (I) denotes an adaptive attacker (cf. Fig. 1), (II) model hardening (passive defense), and (III) an active defense.
@@
As offensive and defensive policies are strictly competitive, we can define a reward $\rho$ of the defensive policy as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$. Then by fixing the optimal evasive policy $\pi_{\theta^{*}}^{\mathcal{A}}$, and updating $\pi_{\phi}^{\mathcal{D}}$ with expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau_{i})]$ gradients, at convergence the resulting ${\pi^{\mathcal{A}}_{\theta^{*}}}$ will be the optimal defensive policy. ∎
@@ III-C Adversarial Markov Games
By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently. As adaptive attacks and defenses are logical consequences of each other, their composition results in a turn-taking competitive game. A precise game-theoretic formulation requires however the exact analytical description of the environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this as well as most other cybersecurity contexts. Model-free methods however can learn optimal offensive and defensive responses directly through interaction with their environment [12, 42], obviating the need to learn a model of it or to find exact solutions to the bi-level optimization problem in Eq. (1) that is NP-hard [43].
To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al. [44] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information. TT-POMGs are an apt formalism for decision-based attacks and defenses, with the additional useful property that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
The competition underlying adversarial example generation has been explored in no-box and white-box settings [27, 45]. We instead focus on decision-based, interactive environments, and assume that they have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior. By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG. The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
The environment that encompasses adversarial attacks, adversarial defenses, and benign queries, we describe it as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2. Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and that consequently the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be also be adaptive. As with plausible assumptions we cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game. When the competing agents that hold beliefs about each other are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]]. In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [46, 47]. In this work we consider more involved recursive reasoning out of scope, as AMGs can be solved by single-agent RL algorithms, and perform the empirical evaluation without building explicit models of opponent behavior.
@@ IV Environment Specification
The empirical study we conduct in Section V reflects diverse instantiations of the general theoretical framework introduced in Section III. When working forward from the theoretical to the practical, concrete design choices have to be made when specifying the latter, choices that can have considerable influence on the results. To elucidate our proposed robustness evaluation methodology, in this section we provide the concrete details on the threat model and the environment.
@@
[Threat Model]. Our AMG framework describes a two-player competitive game; while extensible to more players, in this work we assume that at a given moment only one attack takes place. From the defensive perspective, incoming queries can be either benign or part of an attack. An assumption that influences the effectiveness of stateful detection is that queries can be attributed to UIDs, e.g., an IP address or a user account. However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be required to query the model. To address this, we treat queries irrespective to their source. This is a strictly more challenging setting for stateful defenses, where we operate solely on the content of queries and not on any other metadata, similar to [23]. Unlike Blacklight however, instead of rejecting queries, something that by itself can provide more information to the adversary and thus facilitate evasion (cf. OARS [24], LABEL:tab:comparison), we misdirect by returning the second highest probability class. Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution and so that is not trivial for a defense to tell adversarial noise apart. In summary, the black-box threat model we consider is delineated as follows:
  • Assets: Trained and deployed model $\mathcal{M}$ with corresponding weights ${w}$.
  • Agents: Adversary / Defender / Benign user.
  • Adversary Goal: Generate minimal perturbation adversarial examples in as few queries as possible, while evading the defense.
  • Defender Goal: Stop the adversary from generating adversarial examples, while preserving the correct functionality of the model $\mathcal{M}$ on benign users.
  • Adversary Knowledge: The model $\mathcal{M}$ is known as the black-box function that transforms inputs $x\in[0,1]^{d}$ to outputs $c\in[m]$, $m$ being the number of classes. The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ exists or not.
  • Defender Knowledge: The defender observes only the content of incoming queries, without knowing if they come from a benign user or the adversary.
  • Adversary Capabilities: Adapt the parameters of the attack and optionally any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
  • Defender Capabilities For each query $x$, decide between answering truthfully with the true prediction $C(x)=c_{T}$ or misdirect with the second highest probability class $c_{S}$.
Figure 3: Misdirection in a two-dimensional decision boundary. The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision. $x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
@@
[Similarity]. Decision-based attacks typically follow a policy that generates successive queries: these exhibit degrees of similarity which can be quantified by an appropriate $l_{p}$ norm. If that norm is computed on the original inputs however, an adversary can adapt (see Figure 1) by employing evasive transformations the model is invariant to and bypass the similarity detection. To account for this capability, we train a Siamese network with contrastive loss in order to learn a latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs. For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign or undetected ones.
@@
[Adaptivity]. No evaluation in AML is complete without considering adaptive adversaries; a notion we expand in this work, that is with the instruments to bypass the defense and their optimal configuration. As stateful defenses are so far similarity based, to bypass them intuition points towards input transformations the model is invariant to. For a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))\approx F(x_{t})$. Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold. As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
@@
At this point we should inquire what is the correct composition of transformations $T$ to apply. When shall $T$ be applied, and how does it affect the attack fundamentals? The transformations $T$ can be considered as a set of additional controls, and like attack parameters they themselves can be suboptimal out-of-the-box [15]. Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense. Their trade-off illustrates why the twofold definition of adaptive is necessary in AML evaluations: first to impart the tools to accomplish to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
@@
[Agents & Environments]. Unlike common competitive games, in AMGs the two players have different action and state sets. AMGs are also asymmetric in the playing cadence: while the defender plays every round, the adversary might wait one to several rounds; HSJA for example is controlled on the iteration rather on the query level. Training is complicated further given that the experience upon which each agent learns arrives only after the opponent moves. We address these complications by developing custom learning environments for agents with asymmetries, with delayed experience collection, and asynchronous training, build with the OpenAI Gym and Stable-Baselines3 libraries.
@@
[States & Actions]. For the definition and the rationale behind the states we use, we point the reader to Appendix B. For actions, we control BAGS through 4 parameters: orthogonal step size, source step size, mask bias, and Perlin bias. HSJA is controlled by 3: the gradient estimation radius, the number of estimation queries, and the jump step size. All evaluations start from controlling these attack parameters only; if the active defense proves impossible to defeat, we introduce additional knobs that control the magnitude and probability of transformations on the input, with the goal to evade detection while preserving semantic content and hence the correct classification. The range of transformations we experimented with as well as their magnitude and probability are listed in Table V. Finally, in both BAGS and HSJA the active defense consists of an 1-dimensional continuous action that controls the radius of confinement $a$, as depicted in Fig 3.
@@
[Rewards]. Success in an RL task relies heavily on how it is rewarded. Engineering an effective reward function is non-trivial and hides intricacies, as reward hacking and specification gaming are common phenomena and the learned behavior can vary [48]. For adversaries, the rewards we experimented with are variations on minimizing the distance to the original example – with extra reward shaping based on the fundamental operation of each attack – while defenders are rewarded or penalized for intercepting adversarial or benign queries respectively. The rewards are described in Appendix B.